OpenStudy (anonymous):

@Blank  @jim_thompson5910

OpenStudy (anonymous):

@Blank 

OpenStudy (anonymous):

What do u need help with? :)

OpenStudy (anonymous):

In a random sample of 700 refreshment-dispensing machines, it is found that an average of 8.1 ounces is dispensed with a natural deviation of 0.75 ounce. 28. Find the standard error of the mean. Round your answer to the ten-thousandths place.

OpenStudy (anonymous):

29. Find the probability that the mean of the population will be less than 0.085 ounce from the mean of the sample. Round your answer to the tenths place.

OpenStudy (anonymous):

No. just an answer.

OpenStudy (anonymous):

30. Find the probability that the true mean is between 8.157 and 8.185. Round your answer to the tenths place.

OpenStudy (anonymous):

And thats it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

The answer to number 29 is a percent

OpenStudy (anonymous):

thanks. and 30?

OpenStudy (anonymous):

it says round the answer to the tenths place @jim_thompson5910 can you help?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

which one?

OpenStudy (anonymous):

can you help on #30?

jimthompson5910 (jim_thompson5910):

convert those raw scores to z scores use: z = (x-mu)/sigma

jimthompson5910 (jim_thompson5910):

tell me what z scores you get

OpenStudy (anonymous):

turn 8.157 and 8.185 into z scores right?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

sigma=700 right?

jimthompson5910 (jim_thompson5910):

n = 700 (sample size)

jimthompson5910 (jim_thompson5910):

sigma = 0.75

OpenStudy (anonymous):

z=0.076 and 0.113

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now use this table http://www.math.upenn.edu/~chhays/zscoretable.pdf

jimthompson5910 (jim_thompson5910):

to find the area to the left of z = 0.08 (I rounded z = 0.076 to 2 decimal places)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now what? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what area did you get?

OpenStudy (anonymous):

Wait Im confused

jimthompson5910 (jim_thompson5910):

about what

OpenStudy (anonymous):

From where you said use the table, what did you want me to do?

jimthompson5910 (jim_thompson5910):

the table lets you find the area under the curve to the left of a given z score

jimthompson5910 (jim_thompson5910):

do you remember how to read the table?

OpenStudy (anonymous):

ok but what score do i find?

jimthompson5910 (jim_thompson5910):

z = 0.08 locate the 0.0 row, then find the 0.08 column the value in that row/column combo is the area to the left of z = 0.08

jimthompson5910 (jim_thompson5910):

look on page 2

OpenStudy (anonymous):

0.5319

jimthompson5910 (jim_thompson5910):

what is the area to the left of z = 0.11

jimthompson5910 (jim_thompson5910):

I'm getting z = 0.11 from rounding z = 0.113

OpenStudy (anonymous):

what area? on the graph?

jimthompson5910 (jim_thompson5910):

the table provides the value

jimthompson5910 (jim_thompson5910):

this whole table tells you the area to the left of a given z score

OpenStudy (anonymous):

0.0?

jimthompson5910 (jim_thompson5910):

0.11 starts with 0.1 look in the 0.1 row

OpenStudy (anonymous):

ok then what

jimthompson5910 (jim_thompson5910):

look in the column that finishes up z = 0.11

jimthompson5910 (jim_thompson5910):

0.1 is already taken care of, so the missing bit is that last 1 that's what the column 0.01 means ie 0.11 = 0.1 + 0.01

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0.5438

jimthompson5910 (jim_thompson5910):

now subtract the two areas

OpenStudy (anonymous):

0.5319-0.5438=-0.0119 OR 0.5438-0.5319=0.0119

jimthompson5910 (jim_thompson5910):

the area will be positive, so 0.5438-0.5319=0.0119

jimthompson5910 (jim_thompson5910):

the area between z = 0.08 and z = 0.11 is roughly 0.0119

jimthompson5910 (jim_thompson5910):

that means the area between x = 8.157 and x = 8.185 is also roughly 0.0119

OpenStudy (anonymous):

Ok now what

jimthompson5910 (jim_thompson5910):

the area under the curve is the same as finding the probability (think of randomly throwing a dart onto the right area)

jimthompson5910 (jim_thompson5910):

Find the probability that the true mean is between 8.157 and 8.185. Round your answer to the tenths place. 0.0119 rounds to 0.01 that's roughly a 1% chance

OpenStudy (anonymous):

And thats the answer? 1%?

jimthompson5910 (jim_thompson5910):

oh wait, not 0.01...it should be 0.0 hmm seems odd

jimthompson5910 (jim_thompson5910):

let me recheck

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

oh I used the wrong standard error

jimthompson5910 (jim_thompson5910):

z = (x - mu)/(sigma/sqrt(n)) z = (8.157 - 8.1)/(0.75/sqrt(700)) z = 2.01

jimthompson5910 (jim_thompson5910):

what's the area to the left of z = 2.01 ?

OpenStudy (anonymous):

there is no colomn for 2.01

jimthompson5910 (jim_thompson5910):

combine row 2.0 with column 0.01 notice how they add to 2.01 2.0 + 0.01 = 2.01

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

0.9778

jimthompson5910 (jim_thompson5910):

each row/column combo gives you pretty much every z value from z = -3.49 to z = 3.49

jimthompson5910 (jim_thompson5910):

going in increments of 0.01

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

z = (x - mu)/(sigma/sqrt(n)) z = (8.185 - 8.1)/(0.75/sqrt(700)) z = 2.9985 z = 3.00

jimthompson5910 (jim_thompson5910):

find the area to the left of z = 3.00

OpenStudy (anonymous):

0.9987

jimthompson5910 (jim_thompson5910):

subtract the two

OpenStudy (anonymous):

0.0209

jimthompson5910 (jim_thompson5910):

That still gives 0.0, how annoying honestly, it's better to write it to at least 2 decimal places

OpenStudy (anonymous):

ok i will do that

OpenStudy (anonymous):

Thank you for all of your help. I greatly appreciate it! :)

jimthompson5910 (jim_thompson5910):

you're welcome